Fact: 1 a divided by c = a
Fact: 2 a plus b = a
Fact: 3 abc = 1
If fact 1 and 2 are correct why is fact 3 wrong?
What are the value of each of the variables in fact 1 and 2?

Answer by venugopalramana(3286)

Fact: 1 a divided by c = a
A/C=A....
CASE 1.A=0
CASE 2.IF A IS NOT ZERO THEN C=A/A=1
Fact: 2 a plus b = a
A+B=A.......B=0
Fact: 3 abc = 1
CASE 1.
ABC=0*0*C=0..WHATEVER MAY BE C
HENCE ABC CANNOT EQUAL 1
CASE 2.
IF A IS NOT EQUAL TO ZERO THEN
ABC=A*0*1=0...WHATEVER MAY BE A
HENCE ABC CANNOT EQUAL 1
If fact 1 and 2 are correct why is fact 3 wrong?
What are the value of each of the variables in fact 1 and 2?